OpenStudy (anonymous):

Determine the function f satisfying the given conditions. f'(x)= e^(x/3) f(0) = 4 f(x) = A e^(Bx) + C, b= c=

OpenStudy (anonymous):

i found A= 3

OpenStudy (experimentx):

integrate the f'(x)

OpenStudy (experimentx):

that will be 3e^(x/3) + C

OpenStudy (experimentx):

f(0) = 3e^(0/3) + C 4 = 3 + C or, C = 1, Ans: f(x) = 3e^(x/3) + 1

OpenStudy (anonymous):

b= 1/3?

OpenStudy (anonymous):

or just 3

OpenStudy (experimentx):

f(x) = 3e^(x/3) + 1 is the final answer.

OpenStudy (experimentx):

B would definitely equal to 1/3 by comparison

OpenStudy (anonymous):

thanx